OpenStudy (anonymous):

please help :) change to vertex form using complete the square: 1.) x^2-4x=-1 2.) x^2+2x=-5

jimthompson5910 (jim_thompson5910):

What is the x coefficient for `x^2-4x=-1` ?

OpenStudy (mathmale):

Here you are given two separate quadratic equations. Your job is to rewrite these equations so that they appear in "vertex form," which looks like y = a(x-h)^2 + k, where the vertex is represented by (h,k). Have you seen this before? Secondly, have you used "completing the square" recently? How might I help you?

OpenStudy (anonymous):

hi @mathmale and thank you for responding. I just learned completing the square but i forgot how to use if for these problems. Can you please break it dow for me on the first one and i can try to solve the second myself?

jimthompson5910 (jim_thompson5910):

@study_rocks hopefully you see that the x coefficient of `x^2-4x=-1` is -4? agreed? or no?

OpenStudy (anonymous):

I'm just supposed to turn the problem into vertex form like @mathmale said. But thanks for trying to help me i appreciate it